The three points (5,3), (-2,1),and(-1,-5)are vertices of a triangle. Determine the equations which can be used to model each side of the triangle.
Answer by Alan3354(69443)

The three points (5,3), (-2,1),and(-1,-5)are vertices of a triangle. Determine the equations which can be used to model each side of the triangle.
-------
Label them A(5,3), B(-2,1),and C(-1,-5)
------------
To find the line thru AB:
Find the slope:
slope m = diffy/diffx = (1-3)/(-2-5)
m = 2/7
---
Use y = mx + b and either point to find b, the y-intercept
3 = (2/7)*5 + b
b = 11/7
--> y = (2/7)x + 11/7
Or 2x - 7y = -11/7
That line passes thru A & B
------
Using determinants for BC
---
|+x +y +1|
|-2 +1 +1| = 0
|-1 -5 +1|
----------------
x*(1+5) - y*(-2+1) + 1*(10+1) = 0
6x + y = -11 passes thru B & C
------
Use either method to find AC